40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 353     Entry time: Tue Apr 23 10:21:12 2019
Author: Joe 
Type: Optics 
Category: Configuration 
Subject: Moving the spots to the centre of the curved mirrors 

[Koji,Philip, Liyuan, Joe]

CM1:

We moved the curved mirrors to these positions:

inner = 0.807mm

outer = 0.983 mm

CM2:

inner = 0.92 mm

outer = 0.85 mm

To do this so that realignment was easier, we moved the screws in steps of 5um. We alternated which mirror we adjusted so that we could monitor with a wincam how well aligned the beam into the cavity was. We only moved the cavity mirrors a small amount so we could still see higher order mode flashes transmitted through the cavity (e.g.TM03 modes). We would then improve the input alignment, and then move the cavity mirrors some more. Once the mirrors were adjusted according to http://nodus.ligo.caltech.edu:8080/OMC_Lab/190422_195450/misalignment4.pdf the spot positions looked near the middle of the curved mirrors (using a beam card). We began beam walking but we ran  out of range of the bottom periscope screws in the yaw dof. We tried using the third screw to move the mirrror in both yaw and pitch, hopefully this will let move the mirror such that we can use the just the yaw screw. This screw also ran out of range, so we decided that the cavity needed a small adjustment.

The curved mirrors were moved slightly (>5um) and then we tried to get alignment. By using the fibre coupler translation stage, we move the beam side ways slightly, and then tried to get the periscope mirrors back to a position where the screws could move the mirrors. Once we had an ok alignment, we checked the beam. It looked like it was pretty close to the centre of the curved mirrors, which is where we wanted it to be.

We then tried locking the cavity, although the error signal was quite small. The adjusted the input offset and gain of the servo (there is apparently some problem to do with the input and output offsets). Once the cavity was locked we could make the final adjustments to aligning. We still ran out of range on the periscope. We decided to move the breadboard with the fibre coupler and mode matching lenses on it. Because we knew that the cavity was aligned such that the beam hits the centres of the curved mirrors, we could regain flashes quite quickly. We saw the error signal go down, but eventually this decrease was just to do with the beam clipping on the periscope mirrors. We moved the spot back to where we ok aligned, and slid the periscope so we were not clipping the mirror. This worked very well, and then optimised the alignment.

We then tried to improve the mode matching. 

We took photos of the spot positions (quite near the center) and made the detuned locking measurement. The fitting of the data (attachment 1) wsa 1.1318m (what error should we put here?).

I think the order we did things in was:

  • turning anti clockwise on the fibre coupler and misalign the diode, we measured the modespacing.
  • returned the alignment for the photodiode, and realign fibre couple.
  • miss align the photodiode horizontally, and then used fibre coupler to maximise the peak higher order mode peak height. We then used the PD again to make the peak height bigger.
  •  
Attachment 1: FSR_detuned_locking.pdf  20 kB  Uploaded Tue Apr 23 17:35:26 2019  | Hide | Hide all
FSR_detuned_locking.pdf
Attachment 2: CM1_IMG_7702.JPG  4.384 MB  Uploaded Tue Apr 23 20:21:12 2019  | Hide | Hide all
CM1_IMG_7702.JPG
Attachment 3: CM2_IMG_7704.JPG  4.082 MB  Uploaded Tue Apr 23 20:21:12 2019  | Hide | Hide all
CM2_IMG_7704.JPG
ELOG V3.1.3-